Find the coordinates of point P on AB such that PA/PB=3/4,where A(3,1) and B(-2,5).
Answer by Alan3354(69443)

Find the coordinates of point P on AB such that PA/PB=3/4,where A(3,1) and B(-2,5).
-------------
PA is 3/7 of AB
PB is 4/7 of AB
-----
x from A to B = -5
y from A to B = +4
---
(3/7)*-5 = -15/7
(3/7)*4 = 12/7
P = [A] + [-15/7,12/7]
P is the point (3-(15/7),1+(12/7))
P is (6/7,19/7)